Grade 12th passMechanics

If the particle has a velocity of 5m/s along x-axis at origin, where it will be located at t=6s

Anchal singh
9 Years agoGrade 12th pass

1 Answer

Ritik
9 Years ago
 a=dv/dt
 
We know that a=2t+1
 
Which gives,
 
2t+1=dv/dt
 
Solve this differential equation.
 
∫0t(2t+1)dt=∫5vdv (limit LHS-0 to t, RHS-5 to v)
 
Which gives us the velocity as a function of time.
 
t^2+t=v−5.
 
Put t=6s to get the velocity at t=6sec
 
v=36+6+5
 
Finally, We get v=47m/s